OpenStudy (anonymous):

Let f(t)=t(4-t)^(1/2) on the interval [-1,3] . Find the absolute maximum and absolute minimum of on this interval.

OpenStudy (anonymous):

\[f(t)=t\sqrt{4-t}\]?

OpenStudy (bahrom7893):

take the derivative... set it equal to 0, also check the undefined part..

OpenStudy (bahrom7893):

and the endpoints

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f'(t)=\sqrt{4-t}-\frac{t}{2\sqrt{4-t}}\] \[f'(t)=\frac{2(4-t)-t}{2\sqrt{4-t}}\] \[f'(t)=\frac{8-3t}{2\sqrt{4-t}}\]

OpenStudy (anonymous):

set numerator =0 and solve to find the critical point

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

thank you I really appreciate your help